# Effects of Collaboration on the Performance of Interactive Theme Discovery Systems

### v4

@@ Abstract
11footnotetext: Equal contribution, listed in alphabetical order.
@@ 1 Introduction
Making sense of large amounts of textual data is a common task in many academic disciplines and is generally achieved through qualitative methods such as Thematic Analysis (braun-2006) and Grounded Theory (glaser1968discovery). In these methods, researchers manually search through documents to identify common abstract themes that emerge from the data in a process commonly referred to as "inductive coding". However, manual inductive coding becomes infeasible as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process doi:10.1146/annurev-polisci-090216-023229; IJoC10675.
In the last decade, several NLP solutions have been proposed to support qualitative research methods. These systems offer different ways to systematically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models smith-2018-loop; fang-2023-user, grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset pacheco-2023-interactive; chew2023llmassistedcontentanalysisusing, or providing live recommendations to users as they code the data dai-etal-2023-llm; gao-2024-collabcoder. To ensure the agency of the researchers, each system seeks to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
@@
Generally, tools are evaluated to highlight the strengths and weaknesses of the proposed approach. Examples include evaluating topic coherence with and without human input fang-2023-user and contrasting machine-assisted code-book generation with a manual code-book generation dai-etal-2023-llm. In this paper, we argue that these evaluation strategies are often too narrow and do not account for the different real-world settings in which qualitative researchers employ HitL systems. In particular, we seek to answer the following questions: (1) Does collaboration setting measurably affect the quality of resulting code-books? (2) Do these findings hold across different NLP approaches?
Collaborative Qualitative Analysis (CQA) is a process where researchers work in teams to analyze a dataset to produce a consensus interpretation. Doing so can produce more reliable results, a richer understanding of the data, and insight that combines multiple perspectives cite-key. While there are several frameworks for conducting CQA, in this work we look at two common but contrasting collaboration settings. The first setting is defined by individual coders working separately before jointly consolidating their findings (asynchronous coding). The second setting is defined by coders working in groups and identifying themes through synchronous discussion (synchronous coding). Each setting has different strengths: asynchronous coding is more flexible as teams can work across different regions and time zones, whereas synchronous coding combines diverse perspectives and streamlines the annotation processes.
We test these settings on three different NLP tools for interactive inductive coding by performing experiments in the synchronous and asynchronous contexts. We choose tools that use different underlying NLP methods for this work: (1) the fang-2023-user system, which uses a human-in-the-loop topic modeling solution, (2) the pacheco-2023-interactive system, which models emerging themes with respect to their relationship to other high-level concepts which can be defined by users, and (3) the chew2023llmassistedcontentanalysisusing system, which leverages large language models (LLMs) to judge the adequacy of code definitions, as well as to broadcast them to a large dataset. By way of testing such varied approaches, we achieve a more representative understanding of how different tools perform in different collaboration settings.
To measure outcomes in synchronous vs. asynchronous collaboration, we propose a comprehensive evaluation strategy that looks at different output qualities. Namely, we are interested in measuring the consistency of outputs between synchronous and asynchronous coding, as well as the cohesiveness, and correctness of the resulting themes within each collaboration setting. Finally, we perform a small-scale user study to draw insights from the experience of the coders under each collaboration setting and make design recommendations according to our findings.
In summary, we make two main contributions: (1) We demonstrate how the collaboration setting in which qualitative research takes place affects the outcome of solutions proposed in the NLP literature, by testing three NLP-assisted inductive coding tools in synchronous and asynchronous coding contexts. (2) We propose a comprehensive evaluation strategy that measures different aspects of the outcome of NLP-assisted inductive coding tools, offering insights into the types and quality of themes that they are able to capture. We hope that our findings will inform the development of language technologies that are in line with the qualitative researcher workflows and offer insights for producing useful qualitative research tool evaluations.
@@ 2 Related Work
The overarching goal of the systems we investigate is to partially automate the qualitative coding process either by inducing topics in an interactive, semi-supervised manner fang-2023-user; smith-2018-loop, by learning user-defined themes interactively pacheco-2023-interactive; gao-2023-coaicoder, or by prompting LLMs with natural language definitions of the observed themes chew2023llmassistedcontentanalysisusing; dai-etal-2023-llm. A separate but related line of work exemplified by gao-2024-collabcoder uses LLMs to generate label recommendations as users perform the coding process. While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to annotate large portions of the dataset without extensive supervision.
Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by the Human-Computer Interaction (HCI) literature jiang-2021-serendipity; feuston2021tools; chen-2018-ambiguity. A common finding in the HCI studies on computer-assisted qualitative research is the researchers’ focus on ambiguity. In traditional settings, researchers can independently code a significant portion of the dataset and meaningfully identify areas of ambiguity. However, when working with large datasets and relying on NLP systems, researchers cannot easily uncover all such areas. Solutions have been proposed to either visualize codes (drouhard-aeonium) or rank document disagreement (zade-2018-disagree) regardless of dataset size. Our evaluation differs from these works since we look at different output qualities of the resulting themes and their assignments, combining signal from group overlaps, relationships in the semantic embedding space, as well as manual post-hoc evaluations. This method of evaluation highlights areas where the coders diverge both with each other and with the model, providing another perspective on the ambiguity question.
The motivating factor for our study is the popularization of HitL systems for qualitative coding in the NLP field. With the growing number of new interactive systems comes a need for standardized evaluation across different real-world environments. Previous evaluation methods introduced with new systems have generally been ad hoc, with experiments conducted variously in group settings  choo-utopian; hoque-convisit; smith-2018-loop, on individual participants  rietz-cody, and through platforms such as MTurk  zade-2018-disagree. Our contribution provides a standardized framework for performing experiments in synchronous and asynchronous settings, as well as a set of metrics for evaluating consistency, cohesiveness, and correctness in experimental results.
@@ 3 Interactive Systems
Topic Models. This category includes systems that use some variation of topic modeling to find emerging themes and facilitate document assignment. These systems benefit from the relative speed of the topic model, which allow users to quickly visualize and explore the dataset. Early exploration incorporated visualizations to help users adjust parameters (Chuang2013DocumentEW), while later works implemented refinement operations that allow users to directly edit topic words and remove documents (smith-2018-loop). However, topic modeling systems are limited by their lack of malleability and predictability. Refinement operations mostly edit topic words, which can have limited impact in the final results.
@@
We select the HitL query-driven topic model (QDTM) system introduced by fang-2023-user. The topic model is initialized by providing input queries (i.e., words that represent concepts of interest for the user) which the model uses to generate the initial topics. In our experiments, users begin by iterating through each topic and naming them based on identified themes. Users then use a set of refinement operations to edit the topic model. They can merge and split topics based on topic words, add, remove, or reorder topic words, and remove documents from topics. The next iteration of the model is only produced when the users choose to apply refinements and the prior model is saved, allowing users to return to prior iterations to test different operations. Once satisfied with the state of the topic model, the user downloads the document distribution for that iteration. Details of our experimental settings can be found in App. A.
@@
Relational Approaches. Relational approaches combine vector semantics and structured inference to model relationships between high-level concepts. Instead of treating themes as distributions over words (as topic models do), these frameworks define themes as distributions over generalized concepts. This reflects the inductive coding process, where researchers identify patterns and concepts that are then synthesized into more abstract themes. However, their computational complexity grows with the number of dependencies considered, which hinders their ability to quickly adapt during coding sessions. Further, they rely on users to define informative concepts, making them less suited for inexperienced researchers.
@@
We select the system introduced by pacheco-2023-interactive, which uses a two-stage relational framework. In the first stage, the system automatically partitions the dataset based on semantic similarity. The users explore each partition to identify themes, assign "good" and "bad" example documents for each theme, and input or correct supporting concepts for each example. In the second stage, the system uses the provided examples and concept relations to map the remaining dataset, only leaving documents unmapped if no theme is a sufficiently good match. The assignment procedure follows as a structured inference approach, where dependencies between concepts and themes are explicitly modeled. The unmapped documents are repartitioned as in the first stage and users are prompted to review unmapped partitions again. The process is iterated until all documents are mapped. More details about the framework and our experimental configuration can be found in App. B
@@
Large Language Models. LLMs are ideal candidates for interactive systems, especially for tasks such as qualitative coding where the model can be prompted to produce themes or explanations without ad-hoc training (llms0shotreasoners). They have been used for theme recommendation (gao-2023-coaicoder), for code conflict resolution (gao-2024-collabcoder), and for automated document assignment (deductivecodingwithLLM). However, the flexibility of LLM outputs also leads to hallucinations, which are only partially addressed by prompt engineering. Models further suffer from biases in training which are difficult to identify and impact their ability to produce quality labels or recommendations (chen-2018-ambiguity). Additionally, their massive size is prohibitive when working with large datasets due to the high cost of inference.
@@
We select the framework introduced by chew2023llmassistedcontentanalysisusing. In their protocol, the human coder first manually codes a representative subset of the data and drafts definitions for each code. The LLM is then prompted to label the data sample with the provided definitions. Agreement is calculated between human and model annotations using GWET’s $AC_{1}$ (gwet2008computing). The prompt is then tweaked iteratively to achieve a satisfactory level of agreement, and the best-performing version is used to prompt the model to code the rest of the dataset. Details about our experimental settings can be found in App. D.
@@ 4 Study Design
Dataset. We use a common dataset for all experiments. It was originally released by pacheco-etal-2022-holistic, and it consists of approximately $85,000$ Tweets about COVID-19 vaccines. All tweets in this corpus were posted by users located in the United States, are uniformly distributed between Jan. and Oct. 2021, and contain labels for vaccination stance (e.g. pro-vax, anti-vax) and morality frames (e.g. fairness/cheating and their actor/targets.) roy-etal-2021-identifying.
Participants. We recruited a group of 20 researchers in NLP and Computational Social Science, 6 female and 14 male, between the ages of 20 and 45. This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities. This group covers the range of researchers likely to use interactive coding systems. All participants were either well-versed in qualitative data analysis, or were explicitly trained by senior researchers to perform the task.
Coding Protocol. At the start of each experiment, participants were provided with a demonstration of all the operations in their respective systems. Every system starts with an initial partition of the data, so participants were instructed to read the first 25 samples in each partition, and manually create/name any themes they identified before freely exploring the rest of the dataset and start performing operations to find more themes.
@@
In the relational system experiments, participants were tasked with selecting example documents for each identified theme, as well as determining concept relations for them. Following pacheco-2023-interactive, the supporting concepts considered were vaccination stances and morality frames (e.g., the identified theme “natural immunity” has an “anti-vax” stance, and is tied to the “purity” frame). Once participants were satisfied with their themes and selections, the system automatically coded the rest of the dataset. Unmapped examples were repartitioned and returned to the participants for a second (and last) round of coding.
In the experiments for the LLM-based system, participants produced natural-language definitions for each identified theme and selected a set of good examples for them. We then prompted the LLM with different task-prompt templates to find the best prompt for each set of participant-generated definitions, which was then used to code the rest of the dataset. Details of the templates, as well as the human-model agreement for the best template can be found in Appendix D.
@@
Table 1: Avg. Consistency between Best Theme Matches across Annotator Groups. Statistically significant using a two-sample unpaired t-test with $p<0.05$. ∗∗ Near statistically significant with $p\approx 0.05$.
@@ 5 Evaluation
We use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings. Our evaluation framework is comprised of three dimensions; consistency, cohesiveness & distinctiveness, and correctness, each of which uses metrics that are well-established in the literature NIPS2008_beed1360; 10.5555/3540261.3540416; pacheco-2023-interactive.
@@
Figure 2: Two sets of annotators use a particular HiTL system to find themes. Since the same theme can be named differently by different annotators, we find the best match. In this example, the annotator 1’s theme "VaxIsBad" has been matched with annotator 2’s theme "antivax". After aligning, we calculate the similarity between these two themes using methods like Jaccard Similarity or Centroid Distance.
Consistency. Coders always run the risk of overgeneralizing or missing key themes, resulting in unclear and unsystematic results (cornish-2014-sage-cqa). We propose consistency as one dimension for measuring systematicity. Building on the idea of inter-coder reliability, we define consistency as the measure of annotators eliciting the same themes from the same set of texts (See Fig. 2). Once the dataset has been annotated using an interactive system, there are two obstacles to measuring consistency when using semi-automated annotations. First, themes with similar names may not share the same annotated documents. Similarly, themes with different names may share the same documents. We measure consistency using the overlap of documents between a given pairs of themes. To do this, we calculate the Jaccard similarity across the documents for two themes with different coders. We then select the maximum Jaccard similarity for each (unique) theme, which we consider to be the “best aligned theme”. In Fig. 2, this is shown in the middle where the colored lines indicate the best match obtained by maximum Jaccard overlap. After alignment, we calculate the average of all matched pairs for comparison across settings.

### v5

@@ Abstract
11footnotetext: Equal Contribution
@@ 1 Introduction
Making sense of large textual datasets is a common challenge across academic disciplines and is traditionally addressed through qualitative methods such as Thematic Analysis (Braun and Clarke, 2006) and Grounded Theory (Glaser et al., 1968). These approaches rely on manual inductive coding, in which researchers identify abstract themes by closely reading the data. However, as datasets grow in size, manual coding becomes impractical, motivating the use of Natural Language Processing (NLP) techniques to automate parts of the analysis process Brady (2019); Hilbert et al. (2019).
In recent years, a range of NLP-based systems have been developed to support qualitative research. These systems assist researchers by uncovering latent semantic structures through topic modeling Smith et al. (2018); Fang et al. (2023), clustering documents and propagating limited human annotations across datasets Pacheco et al. (2023); Chew et al. (2023), or offering real-time coding recommendations Dai et al. (2023); Gao et al. (2024). To maintain researcher agency, such systems typically adopt human-in-the-loop (HitL) strategies that balance automation with manual interpretation.
@@
Previous work typically evaluates HiTL qualitative analysis tools in isolation, focusing on specific technical strengths and weaknesses. For example, by comparing topic coherence with and without human input Fang et al. (2023) or contrasting machine-assisted and manual code-book generation Dai et al. (2023). However, qualitative analysis in practice is often collaborative, with teams of researchers jointly coding and interpreting data Flick (2014), and supported by tools built on diverse methodologies that are applied to datasets with vastly differing characteristics Baden et al. (2022). By abstracting away collaboration settings, methodological variation, and dataset diversity, existing evaluations risk misrepresenting how HitL systems operate in real-world settings and how broadly their findings apply. In this work, we seek to answer the following questions: (1) Does the collaboration setting measurably affect the quality of resulting code-books? (2) Do these findings hold across different NLP approaches? (3) How do dataset characteristics influence the outcomes of NLP-assisted inductive coding tools?
We focus on two common but contrasting collaboration settings: asynchronous coding, where individuals code independently before consolidating results, and synchronous coding, where teams identify themes through live discussion. These settings offer complementary strengths, with asynchronous coding supporting flexibility across time and location, and synchronous coding facilitating shared understanding and efficient coordination. To compare outcomes across these settings, we introduce an evaluation framework that measures consistency between synchronous and asynchronous coding, as well as the cohesiveness and correctness of themes produced within each setting.
We evaluate three NLP-assisted inductive coding tools built on distinct methodological foundations: a human-in-the-loop topic modeling system Fang et al. (2023), a concept-driven thematic modeling approach Pacheco et al. (2023), and an LLM-based system for evaluating and propagating code definitions Chew et al. (2023). To assess generalizability across data characteristics, we test all tools and collaboration settings on two markedly different datasets: a corpus of short social media posts and a collection of advertising texts. We further complement our quantitative analysis with a small-scale user study to capture coder experiences and derive design recommendations.
Our contributions are twofold: (1) we demonstrate that the collaboration setting affects the results of NLP-assisted inductive coding tools across diverse methodologies and datasets, and (2) we provide an evaluation strategy that captures multiple dimensions of analysis quality. Together, these findings aim to inform the design and evaluation of language technologies that better align with real-world qualitative research workflows.
@@ 2 Related Work
The overarching goal of the systems we investigate is to partially automate the qualitative coding process either by inducing topics in an interactive, semi-supervised manner Fang et al. (2023); Smith et al. (2018), by learning user-defined themes interactively Pacheco et al. (2023); Gao et al. (2023), or by prompting LLMs with natural language definitions of the observed themes Chew et al. (2023); Dai et al. (2023). A separate but related line of work exemplified by Gao et al. (2024) uses LLMs to generate label recommendations as users perform the coding process. While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to annotate large portions of the dataset without extensive supervision.
Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by the Human-Computer Interaction (HCI) literature Jiang et al. (2021); Feuston and Brubaker (2021); Chen et al. (2018). Prior HCI work shows that coders place particular emphasis on identifying and resolving ambiguity. In traditional settings, this is supported by an independent close reading of the data. However, in large-scale, NLP-assisted analysis, coders have limited visibility into where such ambiguities arise. Solutions have been proposed to either visualize codes (Drouhard et al., 2017) or rank document disagreement (Zade et al., 2018) regardless of dataset size. In contrast, we look at different output qualities of the resulting themes and their assignments, combining signal from group overlaps and relationships in the semantic embedding space, and perform manual post-hoc evaluations. This method of evaluation highlights areas where the coders diverge both with each other and with the model, providing another perspective on the ambiguity question.
Previous evaluation methods introduced with HiTL systems for qualitative coding have generally been ad hoc, with experiments conducted in various group settings  Choo et al. (2013); Hoque and Carenini (2016); Smith et al. (2018), on individual participants  Rietz et al. (2020), and through platforms such as MTurk  Zade et al. (2018). Our contribution provides a standardized framework for performing experiments regardless of the collaboration modality, using a suite of metrics for evaluating consistency, cohesiveness, and correctness in experimental results.
@@ 3.1 Topic Models
This category includes systems that use some variation of topic modeling to find emerging themes and facilitate document assignment. These systems benefit from the relative speed of the topic model, which allow users to quickly visualize and explore the dataset. Early exploration incorporated visualizations to help users adjust parameters (Chuang and McFarland, 2013), while later works implemented refinement operations that allow users to directly edit topic words and remove documents (Smith et al., 2018). However, topic modeling systems are limited by their lack of malleability and predictability. Refinement operations mostly edit topic words, which can have limited impact in the final results.
@@
We select the HitL query-driven topic model (QDTM) system introduced by Fang et al. (2023). The topic model is initialized by providing input queries (i.e., words that represent concepts of interest for the user) which the model uses to generate the initial topics. In our experiments, users begin by iterating through each topic and naming them based on identified themes. Users then use a set of refinement operations to edit the topic model. They can merge and split topics based on topic words, add, remove, or reorder topic words, and remove documents from topics. The next iteration of the model is only produced when the users choose to apply refinements and the prior model is saved, allowing users to return to prior iterations to test different operations. Once satisfied with the state of the topic model, the user downloads the document distribution for that iteration.
To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as Fang et al. (2021), which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$. The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability. The same initial topic model is provided for all experiments.
3.2 Relational Approaches
@@
Relational approaches combine vector semantics and structured inference to model relationships between high-level concepts. Instead of treating themes as distributions over words (as topic models do), these frameworks define themes as distributions over generalized concepts. This reflects the inductive coding process, where researchers identify patterns and concepts that are then synthesized into more abstract themes. However, their computational complexity grows with the number of dependencies considered, which hinders their ability to quickly adapt during coding sessions. Further, they rely on users to define informative concepts, making them less suited for inexperienced researchers.
@@
We select the system introduced by Pacheco et al. (2023), which uses a two-stage relational framework. In the first stage, the system automatically partitions the dataset based on semantic similarity. The users explore each partition to identify themes, assign "good" and "bad" example documents for each theme, and input or correct supporting concepts for each example. In the second stage, the system uses the provided examples and concept relations to map the remaining dataset, only leaving documents unmapped if no theme is a sufficiently good match. The assignment procedure follows as a structured inference approach, where dependencies between concepts and themes are explicitly modeled. The unmapped documents are repartitioned as in the first stage and users are prompted to review unmapped partitions again. The process is iterated until all documents are mapped. More details about the framework and our experimental configuration can be found in App. C
3.3 Large Language Models
@@
LLMs are ideal candidates for interactive systems, especially for tasks such as qualitative coding where the model can be prompted to produce themes or explanations without ad-hoc training (Kojima et al., 2024). They have been used for theme recommendation (Gao et al., 2023), for code conflict resolution (Gao et al., 2024), and for automated document assignment (Xiao et al., 2023). However, the flexibility of LLM outputs also leads to hallucinations, which are only partially addressed by prompt engineering. Models further suffer from biases in training which are difficult to identify and impact their ability to produce quality labels or recommendations (Chen et al., 2018). Additionally, their massive size is prohibitive when working with large datasets due to the high cost of inference.
@@
We select the framework introduced by Chew et al. (2023). In their protocol, the human coder first manually codes a representative subset of the data and drafts definitions for each code. The LLM is then prompted to label the data sample with the provided definitions. Agreement is calculated between human and model annotations using GWET’s $AC_{1}$ (Gwet, 2008). The prompt is then tweaked iteratively to achieve a satisfactory level of agreement, and the best-performing version is used to prompt the model to code the rest of the dataset. For our experiment, we select Llama 3.2 3B-Instruct as the base LLM for automated labeling. Additional details about our experimental settings can be found in App. E.
@@ 4 Study Design
4.1 Datasets
We perform our experiments using two distinct datasets. The first consists of roughly 85,000 tweets about COVID-19 vaccines posted by users located in the US, uniformly distributed between Jan.-Oct. 2021 (Pacheco et al., 2022). The corpus also contains labels for vaccination stance (e.g. pro-vax, anti-vax) and morality frames (e.g. fairness/cheating and their actor/targets) Roy et al. (2021), which are used as auxiliary concepts for the relational model. The second dataset consists of 5,471 climate related English language ads from the Facebook Ad Library Islam et al. (2023). The ads focus on the US, were shown between Jan. 2021-Jan. 2022, and contain labels for climate change stance. The two datasets differ on a number of dimensions such as domain and purpose; a detailed analysis can be found in Appendix A.
4.2 Participants
We recruited a group of 33 researchers in NLP and Computational Social Science, 9 female and 24 male, between the ages of 20 and 45. This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities. This group covers the range of researchers likely to use interactive coding systems. All participants were either well-versed in qualitative data analysis, or were explicitly trained by senior researchers to perform the task. Due to the large number of experiments in our study, some participants took part in multiple experiments. These participants always performed the asynchronous experiment first to prevent external influence and took part in one synchronous experiment at most.
4.3 Coding Protocol
At the start of each experiment, participants were provided with a demonstration of all the operations in their respective systems. Every system starts with an initial partition of the data, so participants were instructed to read the first 25 samples in each partition, and manually create/name any themes they identified before freely exploring the rest of the dataset and start performing operations to find more themes.
@@ 4.3 Coding Protocol
In the relational system experiments, participants were tasked with selecting example documents for each identified theme, as well as determining concept relations for them. Following Pacheco et al. (2023), the supporting concepts considered were vaccination stances and morality frames (e.g., the identified theme “natural immunity” has an “anti-vax” stance, and is tied to the “purity” frame). Once participants were satisfied with their themes and selections, the system automatically coded the rest of the dataset. Unmapped examples were repartitioned and returned to the participants for a second (and last) round of coding.
In the experiments for the LLM-based system, participants produced natural-language definitions for each identified theme and selected a set of good examples for them. We then prompted the LLM with different task-prompt templates to find the best prompt for each set of participant-generated definitions, which was then used to code the rest of the dataset. Details of the templates, as well as the human-model agreement for the best template can be found in Appendix E.
@@
Climate Sync $0.37(0.31)^{*}$ $0.89(0.14)$ $0.43(0.14)$ $0.27(0.17)$ $\bf 0.95(0.05)^{**}$ $\bf 0.43(0.08)$ $0.09(0.07)^{*}$ $\mathbf{0.95(0.04)}$ $\mathbf{0.37(0.07)}$
Async $\mathbf{0.58(0.29)^{*}}$ $\mathbf{0.93(0.12)}$ $\mathbf{0.46(0.16)}$ $\bf 0.33(0.22)$ $0.94(0.09)^{**}$ $0.42(0.08)$ $\mathbf{0.13(0.07)^{*}}$ $0.94(0.06)$ $0.36(0.06)$
Table 1: Avg. Consistency between Best Theme Matches across Coder Groups. Statistically significant using a two-sample unpaired t-test with $p<0.05$. ∗∗ Near statistically significant with $p\approx 0.05$.
@@ 5 Evaluation
We use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when coding in synchronous and asynchronous settings. Our evaluation framework is comprised of three dimensions; consistency, cohesiveness & distinctiveness, and correctness, each of which uses metrics that are well-established in the literature Ben-David and Ackerman (2008); Hoyle et al. (2021); Pacheco et al. (2023).
@@
Figure 2: Two sets of coders use a particular HiTL system to find themes. Since the same theme can be named differently by different coders, we find the best match. In this example, the coder 1’s theme ‘‘VaxIsBad’’ has been matched with coder 2’s theme ‘‘antivax’’. After aligning, we calculate the similarity between these two themes using methods like Jaccard Similarity or Centroid Distance.
5.1 Consistency
Coders risk overgeneralizing or overlooking key themes, leading to unsystematic results (Cornish et al., 2024). We address this by measuring consistency, defined as the extent to which different coders elicit the same themes from the same texts (Fig. 2). In semi-automated coding, consistency is nontrivial to assess: similarly named themes may cover different documents, while differently named themes may overlap substantially. We therefore measure consistency based on document overlap between themes. Specifically, we compute the maximum Jaccard similarity between each theme and all themes produced by another coder, treating this maximum as the theme’s best alignment. Consistency is then calculated as the average similarity across all aligned theme pairs.